Please show work.
Consider the functions:
f(x)=x^2+4
g(x)=x
a Find f(g(x))
b f(g(1))
c find f(2g(x))

Answer by rothauserc(4718)

a) g (x) = ×, then
f (g (x)) = f (x) = x^2 + 4
:
b) f (g (1)) = f (1) = 1 + 4 = 5
:
c) f (2g (x)) = f (2x) = 4x^2 + 4
: